OpenStudy (anonymous):

Solve and Check your solution. 2x^2+2x=12

OpenStudy (anonymous):

start with \[x^2+x=6\] to make your life easier

OpenStudy (anonymous):

then write as \[x^2+x-6=0\] and factor as \[(x+3)(x-2)=0\] finally set each factor equal to 0 and solve

OpenStudy (anonymous):

x=2,-3

OpenStudy (anonymous):

-3,2 correct?@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

yw